gelphielvr:

(algebra 1) arithmetic sequences question in the replies

gelphielvr:

Luigi0210:

Do you know the arithmetic sequence formula?

gelphielvr:

a + (n – 1)^t

Luigi0210:

Well, close, it's actually: \(\large a_n = a_1 +(n-1)*d \) Where \(a_1\) is the first term and d is the difference between the terms. They already have you \(a_1 \) which is \(a_1 = -7\) To find d, just the find the difference between each term: -13 - (-10) = -3 -10 -(-7) = -3 -16 -(-13) = -3 and so on.. so we get d = -3 Just plug those into the formula: \(\large a_n = (-7) +(n-1)(-3) \) \(\large a_n = -7 +(-3n+3 ) \) Giving you: \(\large a_n = -3n -4 \) as your formula.

gelphielvr:

thank you for the explanation luigi - its been weeks since I've done this so this really helped :]